# Conditional entropy and information of quantum processes

### v1

@@ II.4 Outline of the paper
In the latter sections of the paper, we discuss the methods and results in detail. The structure of this paper is as follows. In Section III, we fix our notations and present the preliminaries and background necessary for the discussion of the results. In Section IV, we provide the definition and formula for the conditional entropy of quantum channels and discuss some of its desirable properties. In Section V, we discuss properties and calculations of the conditional entropy of the channel based on relative entropy, called the von Neumann conditional entropy. We derive lower and upper bounds on the von Neumann conditional entropy for arbitrary channels and analytically and numerically calculate the conditional entropy for some of the families of quantum channels of practical interest. We also provide expressions for conditional entropies based on the min-, max-, and geometric Rényi relative entropies. In Section VI, we discuss the connection between the presence of causal influence in bipartite processes and their conditional entropies. We find expression for the conditional entropy for semicausal (nosignalling from nonconditioning input system to conditioning output system) quantum channels in terms of the minimal conditional entropy of the output state. We also provide a lower bound on conditional entropy for $k$-extendible tele-covariant quantum channels. In Section VII, we discuss the operational interpretation of the conditional entropy of quantum channels in terms of asymmetric quantum hypothesis testing for quantum channel discrimination in Stein’s setting and for communication task called conditional quantum channel merging, a variant of quantum state merging protocol. In Section VIII, we study the conditional entropy-based strong subadditivity property for quantum channels and define the conditional mutual information for quantum channels. Analogous to quantum Markov states, we also define quantum Markov channels and study the properties of approximate quantum Markov channels in terms of their recoverability when nonconditioning systems become partially inaccessible or lost. In Section X, we have extended our definition of conditional entropy for bidirectional quantum channels, where two distant parties hold each input and output pair and discuss their properties. In Section IX, we define multipartite mutual information for multipartite quantum channels and derive the expression for tele-covariant quantum channels in terms of their Choi states. We conclude by providing brief remarks and some of the future directions from this work in Section XI.
@@ III Preliminaries
A composite quantum system $AB$ is associated with a separable tensor-product Hilbert space $\mathcal{H}_{AB}:=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, where $\mathcal{H}_{A}$ denotes the Hilbert space of $A$, $|A|:=\dim{(\mathcal{H}_{A})}$, and in general $|A|\leq+\infty$. Going forward we may use $A$ to denote the Hilbert space of system $A$. $\mathrm{L}(A)$, $\mathrm{Pos}(A)$, $\mathrm{St}(A)$ are the set of all linear operators, positive semidefinite operators, and density operators (states), respectively, defined on the Hilbert space $\mathcal{H}_{A}$. A linear operator $\rho_{A}$ denotes $\rho\in\mathrm{L}(A)$. Quantum states are given by density operators that are positive semidefinite operators with unit trace. Pure quantum states are rank-one density operators. A bipartite state $\rho_{AB}$ is called separable between $A:B$ if it can be written as a convex combination of product states, $\rho_{AB}=\sum_{x}p(x)\omega^{x}_{A}\otimes\sigma^{x}_{B}$ for $\sum_{x}p(x)=1$, $p(x)\in[0,1]$ and $\omega^{x}\otimes\sigma^{x}\in{\mathrm{S}t}(AB)$ for all $x$. If a bipartite state $\rho_{AB}$ is not separable, then it is called entangled. Let $\Gamma_{RA^{\prime}}:=\outerproduct{\Gamma}{\Gamma}_{RA^{\prime}}$ denote a maximally entangled operator, i.e., |$\ket{\Gamma}_{RA^{\prime}}:=\sum_{i=0}^{\min\{|A^{\prime}|,|R|\}-1}\ket{i,i}_{RA^{\prime}}$, where {|$\{\ket{i}\}_{i}$ forms an orthonormal set of vectors.
@@
Let $\mathrm{Q}(A^{\prime},A)$ denote the set of all quantum channels with input $A^{\prime}$ and output $A$. The Choi operator $\Gamma^{\mathcal{N}}$ of a completely positive map is $\mathcal{N}_{A^{\prime}\to A}(\Gamma_{RA^{\prime}})$, where $R\simeq A^{\prime}$, i.e., $|R|=|A^{\prime}|$. The Choi state $\Phi^{\mathcal{N}}$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$, when $|A^{\prime}|,|A|<+\infty$, is its normalized Choi operator, i.e., $\Phi^{\mathcal{N}}_{RA}:=\frac{1}{|A|}\Gamma^{\mathcal{N}}_{RA}$. In general, for a linear positive map $\mathcal{M}_{A^{\prime}\to A}$, $\Phi^{\mathcal{M}}_{RA}:=\mathcal{M}(\Phi_{RA^{\prime}})$. Let $\pi_{A}:=\frac{1}{|A|}\mathbbm{1}_{A}$ denote the maximally mixed state for $|A|<+\infty$. $\operatorname{tr}$ denotes the trace operation and $\operatorname{tr}_{A}$ denotes the partial trace of a subsystem $A$ over a composite system. $\log$ denotes the logarithm with base $2$. $\norm{X}_{p}$ for $p\in[1,\infty]$ denotes the Schatten $p$-norm, $\norm{X}_{p}:=\operatorname{tr}[|X|^{p}]^{1/p}$, $|X|=\sqrt{X^{\dagger}X}$ for $X\in\mathrm{L}(A^{\prime})$. The diamond norm of a completely positive map $\mathcal{N}_{A^{\prime}\to A}$ is $\norm{\mathcal{N}}_{\diamond}:=\sup_{\rho\in\mathrm{St}(RA^{\prime})}\norm{\operatorname{id}_{R}\otimes\mathcal{N}(\rho_{RA^{\prime}})}_{1}$, where it suffices to optimize over pure states with $|R|=|A^{\prime}|$. The average energy of a system $A^{\prime}$ in the state $\rho_{A^{\prime}}$ is $\operatorname{tr}[\widehat{H}_{A^{\prime}}\rho_{A^{\prime}}]$, where $\widehat{H}_{A^{\prime}}$ denotes the Hamiltonian of $A^{\prime}$. An isomteric channel $\mathcal{V}_{A^{\prime}\to A}$ is a quantum channel with a single Kraus operator, $\mathcal{V}_{A^{\prime}\to A}(\cdot)=V(\cdot)V^{{\dagger}}$, where $V_{A^{\prime}\to A}$ is an isomtery operator, i.e., $V^{\dagger}V=\mathbbm{1}_{A^{\prime}}$ and $VV^{{\dagger}}=\Pi_{A}$ for some projector $\Pi_{A}$ on the support of $A$. An unitary channel $\mathcal{U}_{A^{\prime}\to A}$ is an isomteric channel with $|A^{\prime}|=|A|$. Let $\mathrm{SC}((A^{\prime},A),(C^{\prime},C))$ denote the set of all superchannels that map $\mathrm{Q}(A^{\prime},A)$ to $\mathrm{Q}(C^{\prime},C)$. The action of $\Theta$ on any quantum channel $\mathcal{N}_{A\rightarrow A^{\prime}}$ can be written as a composition of quantum channel as follows [38]:
@@ III.1 Generalized state divergence
A generalized state divergence $\mathbb{D}(\cdot\|\cdot)$ is a function between two positive semidefinite operators defined on the same Hilbert space that is monotonically nonincreasing under the action of an arbitrary quantum channel. That is, generalized state divergence $\mathbb{D}(\rho\|\sigma)$ for any $\rho\in\mathrm{St}(A^{\prime})$ and $\sigma\in\mathrm{Pos}(A^{\prime})$ and an arbitrary quantum channel $\mathcal{N}_{A^{\prime}\to A}$ satisfies
@@ III.1.1 Examples of generalized state divergence
if $\alpha\leq\beta$, for $\alpha,\beta\in(0,1)\cup(1,\infty)$. For certain range of $\alpha$, the sandwiched Rényi relative entropy $\widetilde{D}_{\alpha}(\rho\|\sigma)$ is a generalized state divergence: Let $\mathcal{N}\in\mathrm{Q}(A^{\prime},A)$ and let $\rho\in\mathrm{St}(A^{\prime})$ and $\sigma\in\mathrm{Pos}(A^{\prime})$, then for all $\alpha\in[1/2,1)\cup(1,\infty)$,
@@
$D_{\max}(\rho\|\sigma)\coloneqq\inf\{\lambda:\ \rho\leq 2^{\lambda}\sigma\}=\log\norm{\sigma^{-1/2}\rho\sigma^{-1/2}}_{\infty},$ (16)
@@ III.2 Generalized channel function and divergence
A generalized channel function $\mathbb{D}[\cdot\|\cdot]$ between any two completely positive maps, quantum channel $\mathcal{N}_{A^{\prime}\to A}$ and a completely positive map $\mathcal{M}_{A^{\prime}\to A}$, can be defined as [31, 23]
@@
where $\mathbb{D}(\cdot\|\cdot)$ is a generalized state divergence. It suffices to restrict $\rho_{RA^{\prime}}$ in (19) to pure states and $R\simeq A^{\prime}$. The above definition is generalization of [31, Definition II.2] (see [23, Definition 3]).
If the input states to the channel are accessible only from a set $\mathsf{Access}(RA^{\prime})\subset\mathrm{St}(RA^{\prime})$ of states because of some physical constraints, then we define the resource-constraint generalized channel function between a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ and a completely positive map $\mathcal{M}_{A^{\prime}\to A}$ as (cf. [23, Eq. (67)])
@@
If we take the supremum over all the states $\rho_{RA}$ in the l.h.s. of the above inequality, then the left-hand side. will become $\mathbb{F}[\mathcal{N}\|\mathcal{M}]$. Let $\psi_{RA}$ be the state realizing the supremum of $\mathbb{F}[\mathcal{N}\|\mathcal{M}]$. Now let us take the supremum over all the states $\rho_{RA}$ in the r.h.s. of the above inequality. Clearly $\psi_{RA}$ is also a valid choice of $\rho_{RA}$ to calculate the supremum of $\mathbb{G}(\cdot\|\cdot)$ and we have the following inequality $\mathbb{F}(\mathrm{id}_{R}\otimes\mathcal{N}_{A\rightarrow A}(\psi_{RA})\|\mathrm{id}_{R}\otimes\mathcal{M}_{A\rightarrow A}(\psi_{RA}))=\mathbb{F}[\mathcal{N}\|\mathcal{M}]\leq\mathbb{G}(\mathrm{id}_{R}\otimes\mathcal{N}_{A\rightarrow A}(\psi_{RA})\|\mathrm{id}_{R}\otimes\mathcal{M}_{A\rightarrow A}(\psi_{RA}))\leq\mathbb{G}[\mathcal{N}\|\mathcal{M}]$, which proves our claim. The similar argument can be used to prove $\mathbb{F}_{\mathrm{A}ccess}[\mathcal{N}\|\mathcal{M}]\leq\mathbb{G}_{\mathrm{A}ccess}[\mathcal{N}\|\mathcal{M}]$. ∎
@@
A bidirectional generalized channel function $\mathbb{D}^{\leftrightarrow}[\cdot\|\cdot]$ between two bipartite completely positive maps $\mathcal{N}_{A^{\prime}B^{\prime}\to AB},\mathcal{M}_{A^{\prime}B^{\prime}\to AB}$ is given by
$\mathbb{D}^{\leftrightarrow}[\mathcal{N}\|\mathcal{M}]:=\sup_{\rho\in\mathrm{St}(R_{A}A),\omega\in\mathrm{St}(R_{B}B^{\prime})}\mathbb{D}(\mathcal{N}(\rho_{R_{A}A^{\prime}}\otimes\omega_{R_{B}B^{\prime}})\|\mathcal{M}(\rho_{R_{A}A^{\prime}}\otimes\omega_{R_{B}B^{\prime}})),$ (24)
@@
Before we proceed further, we make an important observation regarding the monotonicity of the generalized channel function. The proof of the theorem below is directly lifted from [22, 23]. See detailed proof in Appendix B for the sake of completeness.
@@
The generalized channel function $\mathbb{D}\left[\mathcal{N}\|\mathcal{M}\right]$ between an arbitrary pair of completely positive maps, $\mathcal{N}_{A^{\prime}\rightarrow A}$ is a quantum channel and $\mathcal{M}_{A^{\prime}\rightarrow A}$ is a completely positive map, is nonincreasing under the action of a quantum superchannel $\Theta\in SC((A^{\prime},A),(C^{\prime},C))$,i.e.,

### v2

@@ II.4 Outline of the paper
In the latter sections of the paper, we discuss the methods and results in detail. The structure of this paper is as follows. In Section III, we fix our notations and present the preliminaries and background necessary for the discussion of the results. In Section IV, we provide the definition and formula for the conditional entropy of quantum channels and discuss some of its desirable properties. In Section V, we discuss properties and calculations of the conditional entropy of the channel based on relative entropy, called the von Neumann conditional entropy. We derive lower and upper bounds on the von Neumann conditional entropy for arbitrary channels and analytically and numerically calculate the conditional entropy for some relevant quantum channels. We provide expressions for the conditional entropies based on the min-, max-, and geometric Rényi relative entropies. We numerically evaluate the geometric Rényi conditional entropy and the conditional min-entropy of some channels of practical interest. In Section VI, we discuss the connection between the presence of causal influence in bipartite processes and their conditional entropies. We find expression for the conditional entropy for semicausal (nosignaling from nonconditioning input system to conditioning output system) quantum channels in terms of the minimal conditional entropy of the output state. We provide a lower bound on conditional entropy for $k$-extendible tele-covariant quantum channels. In Section VII, we discuss the operational interpretation of the conditional entropy of quantum channels in terms of asymmetric quantum hypothesis testing for quantum channel discrimination in Stein’s setting and for communication task called conditional quantum channel merging, a variant of quantum state merging protocol. In Section VIII, we study the conditional entropy-based strong subadditivity property for quantum channels and define the conditional mutual information for quantum channels. Analogous to quantum Markov states, we define quantum Markov channels and study the properties of approximate quantum Markov channels in terms of their recoverability when nonconditioning systems become partially inaccessible or lost. In Section IX, we define multipartite mutual information for multipartite quantum channels and derive the expression for tele-covariant quantum channels in terms of their Choi states. In Section X, we have extended our definition of conditional entropy for bidirectional quantum channels, where two distant parties hold each input and output pair and discuss their properties. We conclude by providing brief remarks and some of the future directions from this work in Section XI.
@@ III Preliminaries
A composite quantum system $AB$ is associated with a separable tensor-product Hilbert space $\mathcal{H}_{AB}:=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, where $\mathcal{H}_{A}$ denotes the Hilbert space of $A$, $|A|:=\dim{(\mathcal{H}_{A})}$, and in general $|A|\leq+\infty$. Going forward we may use $A$ to denote the Hilbert space of system $A$. $\mathrm{L}(A)$, $\mathrm{Pos}(A)$, $\mathrm{St}(A)$ are the set of all linear bounded operators, positive semidefinite and bounded operators, and density operators (states), respectively, defined on the Hilbert space $A$. A linear bounded operator $\rho_{A}$ denotes $\rho\in\mathrm{L}(A)$. Quantum states are given by density operators that are positive semidefinite operators with unit trace. Pure quantum states are rank-one density operators. A bipartite state $\rho_{AB}$ is called separable between $A:B$ if it can be written as a convex combination of product states, $\rho_{AB}=\sum_{x}p(x)\omega^{x}_{A}\otimes\sigma^{x}_{B}$ for $\sum_{x}p(x)=1$, $p(x)\in[0,1]$ and $\omega^{x}\otimes\sigma^{x}\in{\mathrm{S}t}(AB)$ for all $x$. If a bipartite state $\rho_{AB}$ is not separable, then it is called entangled. Let $\Gamma_{RA^{\prime}}:=\outerproduct{\Gamma}{\Gamma}_{RA^{\prime}}$ denote a maximally entangled operator, i.e., |$\ket{\Gamma}_{RA^{\prime}}:=\sum_{i=0}^{\min\{|A^{\prime}|,|R|\}-1}\ket{i,i}_{RA^{\prime}}$, where {|$\{\ket{i}\}_{i}$ forms an orthonormal set of vectors. We know $\Gamma_{R_{A}R_{B}:A^{\prime}B^{\prime}}=\Gamma_{R_{A}A^{\prime}}\otimes\Gamma_{R_{B}B^{\prime}}$.
@@
Let ${\mathrm{C}P}(A^{\prime},A)$ denote the set of all completely positive maps $\mathcal{M}_{A^{\prime}\to A}$ that outputs bounded operators $\sigma\in{\mathrm{P}os}(A)$ if input is a bounded operator $\rho\in{\mathrm{P}os}(A)$. Let $\mathrm{Q}(A^{\prime},A)$ denote the set of all quantum channels with input system $A^{\prime}$ and output system $A$. A quantum channel $\mathcal{N}_{A^{\prime}\to A}:{\mathrm{S}t}(A^{\prime})\to{\mathrm{S}t}(A)$ is a completely positive, trace-preserving map and $\mathcal{N}_{A^{\prime}\to A}\in{\mathrm{Q}}(A^{\prime},A)$. $\operatorname{id}_{A}$ denotes the identity channel acting on $X\in{\mathrm{L}}(A)$. $\operatorname{id}_{R}\otimes\mathcal{N}_{A^{\prime}\to A}$ may simply be denoted as $\mathcal{N}_{A^{\prime}\to A}$ whenever context is clear. The Choi operator $\Gamma^{\mathcal{N}}$ of a completely positive map is $\mathcal{N}_{A^{\prime}\to A}(\Gamma_{RA^{\prime}})$, meaning $\operatorname{id}_{R}\otimes\mathcal{N}_{A^{\prime}\to A}(\Gamma_{RA^{\prime}})$ where $R\simeq A^{\prime}$, i.e., $|R|=|A^{\prime}|$. The Choi state $\Phi^{\mathcal{N}}$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$, when $|A^{\prime}|,|A|<+\infty$, is its normalized Choi operator, i.e., $\Phi^{\mathcal{N}}_{RA}:=\frac{1}{|A^{\prime}|}\Gamma^{\mathcal{N}}_{RA}$. In general, for a linear positive map $\mathcal{M}_{A^{\prime}\to A}$, $\Phi^{\mathcal{M}}_{RA}:=\mathcal{M}(\Phi_{RA^{\prime}})$. Let $\pi_{A}:=\frac{1}{|A|}\mathbbm{1}_{A}$ denote the maximally mixed state for $|A|<+\infty$. $\operatorname{tr}$ denotes the trace operation and $\operatorname{tr}_{A}$ denotes the partial trace of a subsystem $A$ over a composite system; for $\operatorname{tr}(X_{A})$ to be finite, we need $X\in{\mathrm{L}}(A)$ to be a trace-class operator. $\operatorname{tr}(\mathbbm{1}_{A})=|A|$, which is infinite if $|A|=+\infty$. $\log$ denotes the logarithm with base $2$. $\norm{X}_{p}$ for $p\in[1,\infty]$ denotes the Schatten $p$-norm, $\norm{X}_{p}:=\operatorname{tr}[|X|^{p}]^{1/p}$, $|X|=\sqrt{X^{\dagger}X}$ for $X\in\mathrm{L}(A^{\prime})$. The diamond norm of a completely positive map $\mathcal{N}_{A^{\prime}\to A}$ is $\norm{\mathcal{N}}_{\diamond}:=\sup_{\rho\in\mathrm{St}(RA^{\prime})}\norm{\operatorname{id}_{R}\otimes\mathcal{N}(\rho_{RA^{\prime}})}_{1}$, where it suffices to optimize over pure states with $|R|=|A^{\prime}|$. The average energy of a system $A^{\prime}$ in the state $\rho_{A^{\prime}}$ is $\operatorname{tr}[\widehat{H}_{A^{\prime}}\rho_{A^{\prime}}]$, where $\widehat{H}_{A^{\prime}}$ denotes the Hamiltonian of $A^{\prime}$. An isomteric channel $\mathcal{V}_{A^{\prime}\to A}$ is a quantum channel with a single Kraus operator, $\mathcal{V}_{A^{\prime}\to A}(\cdot)=V(\cdot)V^{{\dagger}}$, where $V_{A^{\prime}\to A}$ is an isomtery operator, i.e., $V^{\dagger}V=\mathbbm{1}_{A^{\prime}}$ and $VV^{{\dagger}}=\Pi_{A}$ for some projector $\Pi_{A}$ on the support of $A$. An unitary channel $\mathcal{U}_{A^{\prime}\to A}$ is an isomteric channel with $|A^{\prime}|=|A|$. Let $\mathrm{SC}((A^{\prime},A),(C^{\prime},C))$ denote the set of all superchannels that map $\mathrm{Q}(A^{\prime},A)$ to $\mathrm{Q}(C^{\prime},C)$. The action of a superchannel $\Theta\in\mathrm{SC}((A^{\prime},A),(C^{\prime},C))$ on any quantum channel $\mathcal{N}_{A\rightarrow A^{\prime}}$ can be written as a composition of quantum channel as follows [41]:
@@ III.1 Generalized state divergence
A generalized state divergence $\mathbb{D}(\cdot\|\cdot)$ is a function between two positive semidefinite operators defined on the same Hilbert space that is monotonically nonincreasing for the states under the action of an arbitrary quantum channel. That is, generalized state divergence $\mathbb{D}(\rho\|\sigma)$ for any $\rho\in\mathrm{St}(A^{\prime})$ and $\sigma\in\mathrm{St}(A^{\prime})$ and an arbitrary quantum channel $\mathcal{N}_{A^{\prime}\to A}$ satisfies
@@ III.1.1 Examples of generalized state divergence
if $\alpha\leq\beta$, for $\alpha,\beta\in(0,1)\cup(1,\infty)$.
For certain range of $\alpha$, the sandwiched Rényi relative entropy $\widetilde{D}_{\alpha}(\rho\|\sigma)$ is a generalized state divergence [27, 29]: Let $\mathcal{N}\in\mathrm{Q}(A^{\prime},A)$ and let $\rho,\sigma\in\mathrm{St}(A^{\prime})$, then for all $\alpha\in[1/2,1)\cup(1,\infty)$,
@@
$D_{\max}(\rho\|\sigma)\coloneqq\log\inf\{\lambda:\ \rho\leq\lambda\sigma\}=\log\norm{\sigma^{-1/2}\rho\sigma^{-1/2}}_{\infty},$ (18)
@@ III.2 Generalized channel function and divergence
A generalized channel function $\mathbb{D}[\cdot\|\cdot]$ between any two completely positive maps, quantum channel $\mathcal{N}_{A^{\prime}\to A}$ and a completely positive map $\mathcal{M}_{A^{\prime}\to A}\in{\mathrm{C}P}(A^{\prime},A)$ can be defined as [34, 24]
@@
where $\mathbb{D}(\cdot\|\cdot)$ is a generalized state divergence. It suffices to restrict $\rho_{RA^{\prime}}$ in (21) to pure states and $R\simeq A^{\prime}$. The above definition is generalization of [34, Definition II.2] (see [24, Definition 3]).
If the input states to the channel are accessible only from a set $\mathsf{Access}(RA^{\prime})\subset\mathrm{St}(RA^{\prime})$ of states because of some physical constraints, then we define the resource-constraint generalized channel function between a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ and a completely positive map $\mathcal{M}\in{\mathrm{C}P}(A^{\prime},A)$ as (cf. [24, Eq. (67)])
@@
If we take the supremum over all the states $\rho_{RA}$ in the l.h.s. of the above inequality, then the l.h.s. will become $\mathbb{F}[\mathcal{N}\|\mathcal{M}]$. Let $\psi_{RA}$ be the state realizing the supremum of $\mathbb{F}[\mathcal{N}\|\mathcal{M}]$. Now let us take the supremum over all the states $\rho_{RA}$ in the r.h.s. of the above inequality. Clearly $\psi_{RA}$ is also a valid choice of $\rho_{RA}$ to calculate the supremum of $\mathbb{G}(\cdot\|\cdot)$ and we have the following inequality $\mathbb{F}(\mathrm{id}_{R}\otimes\mathcal{N}_{A\rightarrow A}(\psi_{RA})\|\mathrm{id}_{R}\otimes\mathcal{M}_{A\rightarrow A}(\psi_{RA}))=\mathbb{F}[\mathcal{N}\|\mathcal{M}]\leq\mathbb{G}(\mathrm{id}_{R}\otimes\mathcal{N}_{A\rightarrow A}(\psi_{RA})\|\mathrm{id}_{R}\otimes\mathcal{M}_{A\rightarrow A}(\psi_{RA}))\leq\mathbb{G}[\mathcal{N}\|\mathcal{M}]$, which proves our claim. The similar argument can be used to prove $\mathbb{F}_{\mathrm{A}ccess}[\mathcal{N}\|\mathcal{M}]\leq\mathbb{G}_{\mathrm{A}ccess}[\mathcal{N}\|\mathcal{M}]$. ∎
@@
A bidirectional generalized channel function $\mathbb{D}^{\leftrightarrow}[\cdot\|\cdot]$ between two bipartite completely positive maps $\mathcal{N},\mathcal{M}\in{\mathrm{C}P}(A^{\prime}B^{\prime},AB)$ is given by
$\mathbb{D}^{\leftrightarrow}[\mathcal{N}\|\mathcal{M}]:=\sup_{\rho\in\mathrm{St}(R_{A}A^{\prime}),\omega\in\mathrm{St}(R_{B}B^{\prime})}\mathbb{D}(\mathcal{N}(\rho_{R_{A}A^{\prime}}\otimes\omega_{R_{B}B^{\prime}})\|\mathcal{M}(\rho_{R_{A}A^{\prime}}\otimes\omega_{R_{B}B^{\prime}})),$ (26)
@@
Before we proceed further, we make an important observation regarding the monotonicity of the generalized channel function. The proof argument of the theorem below is directly lifted from [23, 24]. See detailed proof in Appendix A for the sake of completeness.
@@
The generalized channel function $\mathbb{D}\left[\mathcal{N}\|\mathcal{M}\right]$ between an arbitrary pair of completely positive maps, $\mathcal{N}\in{\mathrm{Q}}(A^{\prime},A)$ and $\mathcal{M}\in{\mathrm{C}P}(A^{\prime},A)$, is nonincreasing under the action of a quantum superchannel $\Theta\in SC((A^{\prime},A),(C^{\prime},C))$, i.e.,